Agenda/Minutes of the VIS Meeting on 2016/9/2
2016/9/2 13:30 - @ Room 210
Participants: Mark, Takahashi, Ohishi, Aso, Hirata, Okutomi, Sato, Tanaka, Ishizaki, Fabian, Shoda, Raffaele
- Top chamber: Installation and leak test will be postponed to January due to delay of cooling test of the cryostat.
- EQ stop, release sensor, dummy payload: Procurement is in the bidding process. Open bidding is on 13th September.
- BF damper: Assembling the LVDT prototype.
- Maraging wire: Stress concentration around the head was pointed out by CRY. Now investigationg.
Type B Past Week Summary - weeks of 8/22 and 8/29
- Fabian, Mark, Zhao-kun and Yokogawa-kun did BS installation at Kamioka:
- Weighed the BF, SF and dummy BS (in mirror box).
- Moved the geophones onto the optical table, clamped them down and took another time series for analysis.
- Aligned the IM, installed the extra trim masses, tested the picos and weighed the whole assembly.
- Placed vacuum flanges and cleaned D-Sub cables ready for routing to the rack.
- Realigned and immobilized the optical table (which had been bumped).
- Made safety improvements on the second floor (new safety rail, new arrangement of floor pieces, caution tape around holes)
- Suspended the dummy BS. Currently however it is not hanging correctly, but is off in the +Y direction by several mm relative to the mirror box.
- We set up the BS OL, using one of the wheeled chests of drawers as a base and using the OL breadboard from the PR2 assembly frame.
- Hirata-san made a 3D assembly model for SR2 and added the optical design.
- Hirata-san worked with Uchiyama-san to solve the 90° rotation problem. It turns out the "box" part with the pistons can be rotated, so we can build the SR2 according to the original design.
- Hirata-san investigated the puzzle of the BF-IM maraging rods, where the 2D drawings and 3D CAD do not agree. Fortunately the actual parts seem to agree with the 3D CAD.
- Hirata-san washed 3 OSEMs (last of initial 10 for BS) and parts for two BFs (for Type A). He moved the BF parts into the clean room and will assemble them with ATC staff from next Wednesday 9/7.
- Type B ongoing issues
Longer cables for BF->PI and SF->PI - RECEIVED
- Longer term to-do list
- Decide position of BS rack - DONE
- Do Inventor assembly for SR2 in chamber and on assembly frame (Hirata) - NEARLY DONE.
- Write more of payload assembly procedure (Fabian) - ONGOING
- Lower breadboard installation to be incorporated into the procedure document.
- Work on BS simulation (Fabian)
- Do BS Simulink model (Mark, Kokeyama)
- Do 2D drawings of SR2 parts (Hirata)
- Redesign SRx IM wire clamps for larger prisms, remachined RMs
- Order SRM/PRM parts (Hirata/Shoda)
- Order SRx fasteners (Hirata)
- Strength test of the stainless steel wire for the bread board may be necessary. Will ask Mirapro if they can do this test (Takahashi)
- Type-Bp modification
- Saito-san made some updates around the LVDT coils. Shoda is implementing them into the 3D CAD model.
- Obuchi-san provided the result of the structure analysis. Shoda will gather the information and write the design document.
- Saito-san is asking the cost estimate to a company and still waiting for the reply. We asked him to ask the cost estimate to another company.
- Shoda is preparing the specification sheets for the ordering.
- We are planing to have the final design reciew from 12 - 14th September.
- Additional ordering for PRs.
- additional masses for IM: Hirata-san made the drawing.
- additional masses for BF: Shoda is making the model.
- IM flags: Hirata-san will make the drawing for the additional machinig.
- TM flags: Discussion is ongoing with Aso-san, Mark, and Hirata-san. (Old design parts will be ordered together with BF modification parts just in case.)
- 1st wire breakers: will be ordered with BF modification parts
- 2nd wire breakers: will be ordered with BF modification parts
- the correts for the hexagon cable clamps: the CAD model of the maraging rod is wrong. Re-design will be done by Hirata-san.
- maraging rods for BFRM: will be ordered with Type-A maraging rods.
- signal cables: will be ordered with Type-A.
- BF EQ stop ring: will be ordered together with the assembly frame.
- Assembly frame
- Design is almost fixed: shown via Seevogh
- Optical bench
- Installation jigs: shown via Seevogh
- Adapters: shown via Seevogh
- wires: shown via Seevogh (the hook analysis is necessary?)
- missing hooks: shown via Seevogh
- Shoda is preparing the specification sheets for the ordering (together with the assembly frame).
- EQ stop: 3sets are in Akeno, un-washed. Takahashi-san will order the washing.
- OB-IM, OB-MD: 3sets are not ordered yet. Takahashi-san will order them.
- Bottom filter cup
The thickness of the top plate of the cup is different from the design. (12mm -> 11mm)
- Mirror preparation
- Hirose-san and Shoda are preparing jigs for the flag detaching. [[|Catoon]]
- Prepare for the final design review. (Shoda)
- Make the design document (Shoda, Fujii)
- Ordering preparation (Shoda)
- Design work around flags (Hirata, Mark)
- Measured location of things in IFI and IMMT chamber with Kasuya-san on 23-24th Aug.
- It smelt bad (musty) outflow air from the vacuum chamber.
- Consider to install IMMT2 suspension from the Top flange of the chamber.
- Discussing with Somiya-san, Hirose-san and Kawamura-san.
- Kitazawa-san (Toyama Univ.) is coming to help coupling measurement of the small OSEMs for IMMT.
- 3D2D drawings of current IMMT suspension almost completed.
- Next week, MCF chamber will be opened.
Watchdog system (Tanaka)
Progress report(For Power circuit re-design)PDF
- Any issue ?
- Insufficient strength of the maraging rods ?
- Asked ATC to do FEA.
- Type-B top wire has the first priority (tell Obuchi-san: Akutsu)
- Will ask KEK to perform the strength test (Takahashi). ~ 3 weeks to complete.
- It is already very likely that at least the top wire for Type-A needs to be re-designed.
- New wire design proposal (Obuchi)
- Need to have a meeting with the manufacturing company to see if they can produce the new design.
- The receptacle part may also need to be modified.
- Replacing it is easy.
- When do we need new wires ?
- Type-Bp: beginning of Dec
- Type-B: middle October
- Type-A: middle January 2017
- Asked ATC to do FEA.
- Local sensor to monitor the TM motion during assembly
Since we will not have the length sensing OpLev during the assembly, we need to prepare another local sensor to measure the transfer function to check the mirrors are suspended properly without touching anywhere.
- Two possibilities
- Use photo sensors to monitor the motion of RM: call Okutomi later (Aso)
- Commercial sensor: only one is available. can be used even with a rough surface like first contact.
Site work plan for the next week
- Mon: BS test hanging (Mark, Fabian, Zhao)
- Tue: BS test hanging (Mark, Fabian, Zhao)
- Wed: BS test hanging (Mark, Fabian, Zhao)
- Thu: BS test hanging (Mark, Fabian, Zhao)
- Travel (Week of 9/5):
- Mark: 5 - 8
- Fabian: 5 - 8
- Travel (Week of 9/12):
- Takahashi: (12-16)
- Shoda: 15-16 (ASJ)
- Mark: 12 - 16
- Fabian: 12 - 16
- Sato: (12-16)
2016/9/9(Fri) 13:30 - in Mitaka